OpenStudy (anonymous):

if it takes 15 min to travel 25km , how long will it take to travel 40 km at the same rate

OpenStudy (anonymous):

24 minutes. \[\frac{ 15 mins }{ 25km }=\frac{ x }{ 40km }\] We want to find the value of x, which represents how long (in minutes) it will take to travel 40 km. So, we need to "cross multiply";

OpenStudy (anonymous):

\[15\times40=600\] \[25\times x=25x\] So, \[600=25x\] Now we need to divide by 25 to get x alone; \[\frac{ 600 }{ 25 }=\frac{ 25x }{ 25 }\] \[24=x\]